If I react 2.9(10)²⁰ molecules of C12H22O11 how many grams of COs will be made?
C12H22O11 + O2 --> CO2 + H2O
1
Expert's answer
2020-12-14T14:36:10-0500

Solution:

The balanced chemical equation:

C12H22O11 + 12O2 --> 12CO2 + 11H2O

According to the equation: n(C12H22O11) = n(CO2) / 12


One mole of any substance contains 6.022×1023 atoms/molecules.

Hence,

n(C12H22O11) = (2.9×1020 molecules / 6.022×1023 molecules) × 1 mol = 0.00048 moles C12H22O11


n(C12H22O11) = n(CO2) / 12

n(CO2) = 12 × n(C12H22O11) = 12 × 0.00048 moles = 0.00576 moles


Moles of CO2 = n(CO2) = Mass of CO2 / Molar mass of CO2

The molar mass of CO2 is 44.01 g mol-1.

Hence,

Mass of CO2 = n(CO2) × M(CO2) = 0.00576 moles × 44.01 g mol-1 = 0.2535 g

Mass of CO2 is 0.2535 grams.


Answer: 0.2535 grams of CO2 will be made.